OpenStudy (mimi_x3):

We know that to determine the EAR which is basically the annual interest rate for compounding frequencies with the adjustments you have to compute the following \[EAR =\left ( 1+\frac{\text{ annual interest rate } }{\text{ compounding periods }} \right ) ^{\text{compounding periods}}-1\] Now when take the limit of shorter and shorter compounding periods meaning continuous compounding we get a formula \( EAR = e^r-1\) Where r is the annual interest rate

OpenStudy (mimi_x3):

@dan815 How did they derive that formula for continuous compounding from the original equation?

OpenStudy (dan815):

cuz

OpenStudy (mimi_x3):

Not sure if i was clear ...-.- so like ask me questions if i didnt make sense

OpenStudy (dan815):

write a differential eqn

OpenStudy (dan815):

wait ill brb.. u try to clarify the question as much as u can

OpenStudy (dan815):

ok bacck

OpenStudy (mimi_x3):

Basically as the compounding periods increase ... lets say from quaterly to annually to monthly and then to daily ... the EAR becomes greater and greater ... and it kinda converges at compounding period of 365 Basically the second formula finds the value it converges to

OpenStudy (dan815):

ya okay

OpenStudy (dan815):

basically umm the normal compounding formula is Total=P*(1+r)^n

OpenStudy (dan815):

then we when u got annual rate being compounded nominally over a period of tie

OpenStudy (mimi_x3):

Ya im trying to find the total interest rate after adjustments meaning with compunding periods

OpenStudy (dan815):

Total=P*(1+r/k)^(n*k)

OpenStudy (mimi_x3):

yes

OpenStudy (dan815):

now u wanna see what happens as k goes to infinite

OpenStudy (mimi_x3):

when k goes infinite

OpenStudy (mimi_x3):

and not exactly with this formula

OpenStudy (mimi_x3):

but with the formula discussing the interest accrued over the year with compounding periods

OpenStudy (dan815):

hmm im thinkg about like

OpenStudy (mimi_x3):

k when it says the interest is 8% compounding quaterly ... that basically more than 8% interest over the yr

OpenStudy (dan815):

binomial series expansiosn

OpenStudy (dan815):

but i dont remember getting growth rates like that

OpenStudy (dan815):

like another way to think about it... okay look you had a continous growth rate

OpenStudy (dan815):

and u got some rate after 1 year

OpenStudy (dan815):

we can write a DE for that

OpenStudy (mimi_x3):

BTWWWW

OpenStudy (mimi_x3):

PLS READ MY FORMULA

OpenStudy (dan815):

ya i se theres -1

OpenStudy (mimi_x3):

No n meaning number of years isnt included

OpenStudy (dan815):

oh okay u wanan see for 1 year rightq

OpenStudy (mimi_x3):

its just abt the interest and compounding periods

OpenStudy (mimi_x3):

yes

OpenStudy (dan815):

ok!

OpenStudy (mimi_x3):

hahaha :P

OpenStudy (dan815):

lemme ask uthis

OpenStudy (dan815):

after 1 year, like after the total compounding period in 1 yeasr

OpenStudy (dan815):

would you expect A=p(1+r)

OpenStudy (dan815):

like lets say its 12% being compounded continously

OpenStudy (mimi_x3):

Nope

OpenStudy (dan815):

okay i see

OpenStudy (dan815):

what wud u expect in 1 year?

OpenStudy (mimi_x3):

hmmmm lemme take a pic of my work one sec

OpenStudy (mimi_x3):

http://puu.sh/fFP73/a353d27252.jpg

OpenStudy (dan815):

okay i see

OpenStudy (mimi_x3):

So basically i showed u that the more its compounded then the more the interest is per year

OpenStudy (mimi_x3):

but then it converges at daily obv

OpenStudy (mimi_x3):

And its showing that u cld use the formula e^r-1 to find the convergence

OpenStudy (dan815):